OpenStudy (anonymous):

Find the real number solutions: 4w^4+40w^2-44=0 I need the steps! Thanks!

OpenStudy (turingtest):

4w^4+40w^2-44=0 4(w^4+10w^2-11)=0 4(w^2-1)(w^2+11)=0 since the second set of parentheses give an imaginary answer, the only one to worry about it the first set, which yields w={-1,1}

OpenStudy (anonymous):

Thanks! :)